OpenStudy (anonymous):

integral (x^2-2x+1)/x dx

OpenStudy (earthcitizen):

\[\int\limits (x^2-2 x+1)/x dx = 1/2 (x-4) x+\ln(x)+c\]

OpenStudy (anonymous):

Can you please show me how you go that

OpenStudy (mimi_x3):

\[\int\limits\frac{x^{2}-2x+1}{x} dx\] => \[\int\limits\frac{x^{2}}{x} -\frac{2x}{x} +\frac{1}{x} dx\] => \[\int\limits x-2+\frac{1}{x} dx\] The integrate it..

OpenStudy (mimi_x3):

Then*

OpenStudy (anonymous):